Hello, just want to make sure I showed the steps right:
3x + 5 = 2x - 9
This is what I did:
5-5 = 0
9-5 = -14
2x - 2x = x
2x - 2x = 0 x = -14 final answer Thank you!!!!!!

Answer by tutorcecilia(2152)

Correct!! I got the same answer, but I used different steps:
3x + 5 = 2x - 9
3x+5-5=2x-9-5
3x=2x-14
3x-2x=2x-2x-14
x=-14